Cholesterol levels for women aged 20 to 34 follow an approximately normal distribution with mean 185 milligrams per deciliter (mg/dl). Women with cholesterol levels above 220 mg/dl are con- sidered to have high cholesterol and about 18.5% of women fall into this category. What is the standard deviation of the distribution of cholesterol levels for women aged 20 to 34?
Answer #1

Given,

= 185 , = ?

We convert this to standard normal as

P(X < x) = P(Z < ( x - ) / )

We have given,

P(X > 220) = 0.185

That is

P(X < 220) = 1 - 0.185

P(X < 220) = 0.815

P(Z < ( x - ) / ) = 0.815

P(Z < (220 -185) / ) = 0.815

From Z table, z-score for the probability of 0.815 is 0.8965

( 220 - 185) / = 0.8965

Solve for

= 39.04